# ViMo: A Generative Visual GUI World Model for App Agent

### v1

@@ Appendix D Evaluation Details
For the results presented in Table 2, we prompt LLM for the instructional accuracy score $s_{ia}$ and action readiness score $s_{ar}$, as shown in Subsection G.7 and Subsection G.8 respectively. A generation is considered successful if “success” appears under “Status” for $s_{ia}$ and “yes” under “ready for action” for $s_{ar}$.
For the results presented in Table 3, we employ an LLM as a judge to assess the alignment between the agent’s simulated actions and the ground truth actions within a given trajectory. Specifically, an agent is prompted to generate the action instructions based on the given GUI and the user goal (see the prompt in Subsection G.9) and the LLM evaluates whether the simulated action leads to the same outcome as the ground truth action (see the prompt in Subsection G.10), a “yes” of the “Status” is calculated as a match.
@@
In Table 4, we categorise apps based on their primary functions into three groups: Leisure, Work, and System. The Leisure category includes apps commonly used for relaxation and entertainment, such as Decathlon, eBay, Flipkart, Amazon, Adidas, Kitchen Stories, Booking.com, YouTube, and Vimeo. The Work category comprises apps typically associated with professional or productivity-related activities, including Gmail, Drive, and Chrome. Lastly, the System category encompasses apps pre-installed in the Android operating system, such as com.android.contacts, com.google.android.dialer, com.google.android.googlequicksearchbox, com.android.settings, com.google.android.apps.maps, and com.android.vending.
Appendix E Additional Experimental Results
@@
Table 9: Evaluation on randomness by running the experiment 3 times (r1-r3) on our sampled test split. “All” denotes the evaluation of the full test split. $s_{gc}$,$s_{ia}$ and $s_{ar}$ are the metrics same with Table 2. $s_{h}$ denotes their harmonic score.
@@
ALL 0.7308 76.07 78.54 0.7583
Table 8 presents additional results of our ViMo applied to APP-Agent and Mobile-Agent-V2. These results further confirm the effectiveness of our ViMo compared to other world models.
@@
Moreover, our method involves random factors, particularly from the use of LLMs. To evaluate their influence, we conducted the experiment three times, as summarised in Table 9 (r1-r3). The results demonstrate that the randomness does not significantly impact the performance or consistency of our method. Additionally, we focused on a randomly selected subset of examples for evaluation, with results from the full test set also included to illustrate that the observed differences are minor, as shown in Table 9 (compare ALL to r1-r3). We consider the subset results to provide an accurate and reliable approximation for our analysis.
Appendix F Additioanl Visualisation
Diverse visualisations are presented in Fig. 8 and Fig. 9. These examples illustrate how our method effectively generates the next GUI based on the given action and current GUI observation, showcasing its ability to produce visually coherent and contextually accurate GUI simulation.
@@ Appendix F Additioanl Visualisation
Moreover, Fig. 10 showcases results generated from a single current GUI paired with different actions, further highlighting the versatility of our approach. Fig. 11 illustrates failure cases where text symbols are not represented as our rectangle-shaped placeholders with a black border and white fill, making them unrecognizable as text symbols. Improving the representation of text symbols remains a potential direction for future work.
@@ G.8 Prompt for action readiness accuracy score ($s_{ar}$)
You are an expert in evaluating the performance of a mobile emulator. The mobile emulator is designed to navigate the UI change based on human instruction. Inputs: UI Screenshot: The mobile emulator generated UI indicating the state of the cellphone’s user interface. User intent: The user’s goal to achieve. Next action: the action will be applied to this UI. Your goal is to determine whether the next action is validated on the UI Screenshot. Please also indicate if it is still in the right App according to the goal. *IMPORTANT* Format your response into a JSON map as shown below: { "Thoughts": <your thoughts and reasoning process>, "In the right App": "yes" or "no" "ready for action": "yes" or "yes", }
G.9 Prompt to generate the action instruction based on the given GUI and the user goal
@@
G.10 Prompt to evaluates whether the simulated action leads to the same outcome as the ground truth action

### v2

@@ A.4 Evaluation
For the results under automatic metrics presented in Table 1, we prompt LLM for the instructional accuracy score $s_{ia}$ and action readiness score $s_{ar}$, as shown in Subsection G.7 and Subsection G.8 respectively. A generation is considered successful if "success" appears under "Status" for $s_{ia}$ and "yes" under "ready for action" for $s_{ar}$. For the user study, we collected 80 generated samples—20 from each of the four world models. We then asked 70 participants to answer three questions on each sample designed to reflect the $s_{ia}$, $s_{gc}$ and $s_{ar}$ scores, as detailed in Subsection G.9. For the $s_{gc}$, participants are asked to rate on a scale from 1 to 5. These scores were then normalised to the [0,1] range in Table 1.
World Model Enhanced App Agent.
@@
In Table 2, we categorised APPs based on their primary functions into three groups: Leisure, Work, and System. The Leisure category includes APPs commonly used for relaxation and entertainment, such as Decathlon, eBay, Flipkart, Amazon, Adidas, Kitchen Stories, Booking.com, YouTube, and Vimeo. The Work category comprises APPs typically associated with professional or productivity-related activities, including Gmail, Drive, and Chrome. Lastly, the System category encompasses APPs pre-installed in the Android operating system, such as com.android.contacts, com.google.android.dialer, com.google.android.googlequicksearchbox, com.android.settings, com.google.android.APPs.maps, and com.android.vending.
Ablation on Iteration Numbers.
ViMo predicts future GUI observations, which can be recursively fed back as input to simulate further into the future. Taking the generative GUI as the current GUI, an agent was prompted to generate the action instructions based on the user goal (see the prompt in Subsection G.10). Then the action instruction and the GUI were fed into ViMo to generate the next GUI. In this study, we defined the iteration number as the number of times ViMo was called. We only use the final output as the signals during the candidate action selection phase, guiding the final selection among potential actions.
Appendix B Additional Experimental Results
Comparison with World Models.
Table 5 compares our ViMo with existing world models under M3A App agent. To further highlight our superiority, Table 9 presents additional results of ViMo applied to T3A, APP-Agent, and Mobile-Agent-V2. One can see from the results that our method consistently outperforms existing world models, clearly demonstrating its superiority.
Table 10: Trajectory synthesis evaluation. "T+L" denotes the accuracy of the whole trajectory with length L.
       World Model        T+1        T+2        T+3        T+4
       w/o world model        22.81        14.04        7.02        0
       Change-text        52.63        26.32        10.53        5.26
       HTML-text        38.60        14.04        12.28        7.02
       HTML-vision        43.86        19.30        10.53        10.53
       UI-diffuser        52.63        29.82        12.28        5.26
       IP2P*        56.14        21.05        10.53        7.02
       ViMo (Ours)        57.89        36.84        14.03        12.28
@@ Appendix B Additional Experimental Results
Table 11: Evaluation on randomness by running the experiment 3 times (r1-r3) on our sampled test split. "All" denotes the evaluation of the full test split. $s_{gc}$,$s_{ia}$ and $s_{ar}$ are the metrics same with Table 1. $s_{h}$ denotes their harmonic score. STD denote the standard deviation from r1 to r3.
@@
STD 0.0057 0.23 0.42 0.0025
ALL $0.7389$ $75.37$ $78.20$ $0.7578$
Generation Error Analysis. As discussed in Subsec. 4.5, our method ViMo can iteratively generate future GUIs. However, as the number of iterations increases, the accumulated error also grows. In addition to the evidence presented in Table 7, we conduct further experiments to analyse this iteration error and compare our approach with existing world models.
To this end, we design a trajectory synthesis evaluation to assess how well the GUIs generated by ViMo align with those observations in real-world environments over longer iterations. In this setup, the generated GUI is leveraged as the input to an App agent to generate the subsequent action, with higher-quality trajectories indicating a GUI more aligned with the real-world environment. Specifically, the GUIs generated by ViMo serve as the observation input for the App agent, which generates actions aimed at achieving the user’s goal. These output actions are then evaluated to reflect whether the GUI representations offer concrete and reliable information for action prediction. This process is repeated for L steps, and we calculate the success rate of the entire L-step trajectory.
We employ an LLM as a judge to assess the alignment between the agent’s simulated actions and the ground truth actions within a given trajectory. Specifically, an agent was prompted to generate the action instructions based on the given GUI and the user goal (see the prompt in Subsection G.10) and the LLM evaluated whether the simulated action lead to the same outcome as the ground truth action (see the prompt in Subsection G.11), a "yes" of the "Status" is calculated as a match.
As shown in Table 10, we compared ViMo against both visual- and language-based world models and demonstrated that while performance decreases across all world models with more iterations, our model significantly outperformed the other methods by providing more accurate and reliable information. This was reflected in higher trajectory prediction accuracy, underscoring the ability of our model to generate GUIs that aligned with the real-world environment.
Randomness Study and Evaluation on Full Test Split.
@@
ViMo involves random factors, particularly from the use of LLMs. To evaluate their influence, we conducted the experiment three times, as summarised in Table 11 (r1-r3). The results demonstrate that the randomness does not significantly impact the performance or consistency of our method. Additionally, we focused on a randomly selected subset of examples for evaluation, with results from the full test set also included to illustrate that the observed differences are minor, as shown in Table 11 (compare ALL to r1-r3). We consider the subset results to provide an accurate and reliable approximation for our analysis.
Figure 4: Qualitative ablation studies. Left: Static text generation. (a) Generating static text via an LLM; (b) Preserving the original text in the image by rendering it as image pixels. Right: STR generation under two input formats—(a) action command and (b) action instruction.
Qualitative Ablation Analysis.
In addition to the quantitative ablation results presented in Table 6, we also provide qualitative comparisons. Fig. 4 (left) illustrates the challenges faced by the LLM in predicting static text under complex spatial layouts. Fig. 4 (right) displays the STR generation of the same user intent but with different action types. It demonstrated that models learned with action commands failed to predict STR that aligns with the user’s intent, whereas action instructions offered a more concrete description, enabling the model to better capture the intent.
Qualitative Generalisation Study.
We studied the generalisation of ViMo in Fig. 5 by providing user actions that were not typically encountered within the App’s standard context. For example, in the Clock App, a user action to "add a file" generated a Drive-style file selection window while retaining the Clock interface. Similarly, in the Kitchen Store App, ViMo can generate content corresponding to the action. These results emphasised ViMo’s generalisation ability facing novel combinations of App observations and user actions.
Appendix C Practical Deployment
In this section, we report the computational efficiency of our method to demonstrate its practicality in real-world applications. The minimum hardware requirement is a GPU with 16 GB of memory. On a V100 GPU, STR image generation takes approximately 8 seconds, and GUI-text prediction takes around 30 seconds. In our setup, the total inference time per request is about 2 minutes, including model loading and communication overhead. Each request involves predicting future GUIs for three different user actions. Table 12 compares the inference time and step accuracy of ViMo with other world models. With an additional 2 minutes of inference time, ViMo achieves a notable accuracy improvement of 6.07%.
Table 12: Inference time and step accuracy comparison across models.
Model Inference Time Step Accuracy (%)
Baseline (T3A) $\sim$4 minutes 43.13
Change-text $\sim$5 seconds 46.64
IP2P* $\sim$1.5 minutes 47.92
ViMo (Ours) $\sim$2 minutes 49.20
Figure 6: False examples where the text symbols are incorrectly represented, making them unrecognizable to indicate the location of text.
Figure 7: Visualisation of ViMo in generating the next GUI. For each example, the action is displayed at the top, with the current GUI shown on the left and the generated GUI on the right.
Figure 8: Visualisation of ViMo in generating the next GUI. For each example, the action is displayed at the top, with the current GUI shown on the left and the generated GUI on the right.
Figure 9: Visualisation of ViMo in generating GUIs given a single current GUI paired with different actions.
Appendix D Limitation
@@
Fig. 6 illustrates failure cases where text symbols are not represented as our rectangle-shaped placeholders with a black border and white fill, making them unrecognisable as text symbols. Improving the representation of text symbols remains a potential direction for future work.
Appendix E Additional Visualisation
Diverse visualisations are presented in Fig. 7 and Fig. 8. These examples illustrate how our method effectively generates the next GUI based on the given action and current GUI observation, showcasing its ability to produce visually coherent and contextually accurate GUI simulations. Moreover, Fig. 9 showcases results generated from a single current GUI paired with different actions, further highlighting the versatility of our approach.
Appendix F Social Impact and Safeguards
Our work, which focuses on predicting future GUI states, is aligned with general advances in the field of machine learning. As such, it carries potential societal implications similar to those associated with generative technologies, such as deepfakes [51], including concerns around misuse, misinformation, or user manipulation. To prevent the high risk of misuse of the proposed method, an additional user commitment will be required for accessing the checkpoint in our forthcoming open release, through which we hope to alleviate the potential misuse while benefiting further research.
@@ G.8 Prompt for action readiness accuracy score ($s_{ar}$)
You are an expert in evaluating the performance of a mobile emulator. The mobile emulator is designed to navigate the UI change based on human instruction. Inputs: UI Screenshot: The mobile emulator generated UI indicating the state of the cellphone’s user interface. User intent: The user goal to achieve. Next action: the action will be applied to this UI. Your goal is to determine whether the next action is validated on the UI Screenshot. Please also indicate if it is still in the right App according to the goal. *IMPORTANT* Format your response into a JSON map as shown below: { "Thoughts": <your thoughts and reasoning process>, "In the right App": "yes" or "no" "ready for action": "yes" or "yes", }
Figure 10: Screenshot of user study example.
G.9 Instructions for User Study
The following prompt provides the instructions for the user study. An example screenshot is shown in Fig. 10.
Question 1: You are presented with a mobile screen (Current GUI). Your task is to evaluate whether the generated GUI correctly reflects the result of the user action applied on the current GUI. Answer "Yes" or "No" to each sample. Question 2: You are presented with Ground Truth Next GUI, the correct screens that should appear after the user performs the given action. In this task, you will evaluate the visual similarity between the Ground Truth Next GUI and AI-generated GUI, scoring from 1-5. Question 3: In this task, you will validate whether a specific user action is valid on the generated GUI. An action is considered valid if the required GUI element is visually presented on the screen. Answer "Yes" or "No" to each sample.
G.10 Prompt to generate the action instruction based on the given GUI and the user goal
@@
G.11 Prompt to evaluates whether the simulated action leads to the same outcome as the ground truth action